# Complete $k$-partite entanglement measure

### v3

@@ 4.2 $k$-PEMo from the minimal sum
[…]
@@
$\quad E^{\prime}_{(k)}(|\psi\rangle^{X_{1}X_{2}X_{3}})\geqslant\frac{1}{2}\sum_{i}p_{i}\left[h(\rho^{X_{1(i)}})+h(\rho^{X_{22(i)}})+h(\rho^{X_{3(i)}})\right]\geqslant E^{\prime}_{(k)}(\rho^{X_{12}X_{2}X_{3}})$
@@
It is clear that the completeness of $E^{\prime}_{(2)}$ is reduced to the subadditivity of the reduced function. This completes the proof. ∎
@@
$\qquad\quad E^{\prime}_{(3)}(|\psi\rangle^{ABCD}|\psi\rangle^{EF}|\psi\rangle^{GHI})=\frac{1}{2}\left[h(\rho^{AB})+h(\rho^{CD})+h(\rho^{GH})+h(\rho^{I})\right],$
@@
$\qquad\quad C_{(3)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})=\frac{1}{5}\left[h(\rho^{AB})+h(\rho^{C})+h(\rho^{FG})+h(\rho^{H})\right],$
@@
$\qquad\quad C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\frac{1}{4}(h_{AB}+h_{C}+h_{DE}+h_{F})=\frac{1}{4}(h_{DE}+h_{F}).$
Then $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{DEF})=\frac{1}{3}(h_{AB}+h_{DE}+h_{F})=\frac{1}{3}(h_{DE}+h_{F})$, i.e., $C_{(3)}$ is not coarsening monotonic. In addition $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}+C_{(3)}(|\psi\rangle^{DEF})=C_{(3)}(|\psi\rangle^{DEF})=\frac{1}{2}(h_{DE}+h_{F})$, so $C_{(3)}$ is not additive. For $|\psi\rangle^{AB}|\psi\rangle^{CD}$, $C_{(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\frac{1}{4}(h_{A}+h_{B}+h_{C}+h_{D})=\frac{1}{2}(h_{A}+h_{C})$ may be not larger than $C_{(2)}(AB|C|D)=2h_{C}/3$.
@@
$\qquad C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{h_{F}}/2<C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{DEF})=\sqrt{2h_{F}/3}$
@@
$C_{G,q(2)}(|\psi_{2}\rangle)=\sqrt{2h_{A}}/2<C_{G,q(2)}(|\psi\rangle^{AB}|\psi\rangle^{D})$
$=$ $C_{G,q(2)}(A|B|CD)=\sqrt{2h_{A}/3},$
i.e., it is neither coarsening monotonic nor tightly coarsening monotonic. $C_{G,\alpha(k)}$ has the same property as that of $C_{G,q(k)}$.
@@ 5 Examples
[…]
@@ 6 Conclusion
We have established the axiomatic postulates for complete measure of the $k$-partite entanglement and presented two classes of $k$-partite entanglement measures. Comparing with the axiomatic postulates of complete $k$-entanglement measure in Ref. [2], both the unification condition and the hierarchy condition were modified accordingly in order to make them in consistence with the type of entanglement considered. Together with the complete $k$-entanglement measure, we get a further progress in characterizing of multipartite entanglement. In comparison, although the $k$-PEM is far different from the $k$-entanglement measure, it has some similarities to the $k$-entanglement measure: all of them can be defined by the reduced function and in such a sense the completeness is always related to the subadditivity of the reduced function. In addition, we can discuss the monogamy and the complete monogamy relations of the $k$-entanglement measure, but it seems not compatible for $k$-PEM. Going further, our result is applicable for other $k$-partite measure of quantum correlations since it is based on the coarsening relation of partitions.
This work is supported by the National Natural Science Foundation of China under Grant Nos. 12471434 and 11971277, the Program for Young Talents of Science and Technology in Universities of Inner Mongolia Autonomous Region under Grant No. NJYT25010, and the High-Level Talent Research Start-up Fund of Inner Mongolia University under Grant No. 10000-2311210/049.

### v4

@@ 4.2 $k$-PEMo from the minimal sum
where $X_{1(i)}|X_{2(i)}|\cdots|X_{m(i)}$ corresponds to $\gamma_{i}^{f}$, $\rho^{X}={\rm Tr}_{\overline{X}}|\psi\rangle\langle\psi|$, and $\overline{X}$ denotes the subsystem complementary to those of $X$. The coefficient “1/2” is fixed by the unification condition when the measures defined via $\mathcal{P}_{k}^{\gamma_{i}^{f}}$ are regarded as unified $k$-PEMs. We define
@@
$E^{\prime}_{(k)}(|\psi\rangle^{X_{1}X_{2}X_{3}})$
$\geqslant$ $\frac{1}{2}\sum_{i}p_{i}\left[h(\rho^{X_{1(i)}})+h(\rho^{X_{22(i)}})+h(\rho^{X_{3(i)}})\right]\geqslant E^{\prime}_{(k)}(\rho^{X_{12}X_{2}X_{3}})$
@@
It is clear that the completeness of $E^{\prime}_{(2)}$ is reduced to the subadditivity of the reduced function. This completes the proof.
@@
$E^{\prime}_{(3)}(|\psi\rangle^{ABCD}|\psi\rangle^{EF}|\psi\rangle^{GHI})=\frac{1}{2}\left[h(\rho^{AB})+h(\rho^{CD})+h(\rho^{GH})+h(\rho^{I})\right],$
@@
$C_{(3)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})=\frac{1}{5}\left[h(\rho^{AB})+h(\rho^{C})+h(\rho^{FG})+h(\rho^{H})\right],$
@@
$C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\frac{1}{4}(h_{AB}+h_{C}+h_{DE}+h_{F})=\frac{1}{4}(h_{DE}+h_{F}).$
Then $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{DEF})=\frac{1}{3}(h_{AB}+h_{DE}+h_{F})=\frac{1}{3}(h_{DE}+h_{F})$, i.e., $C_{(3)}$ is not coarsening monotonic. In addition $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{C})+C_{(3)}(|\psi\rangle^{DEF})=C_{(3)}(|\psi\rangle^{DEF})=\frac{1}{2}(h_{DE}+h_{F})$, so $C_{(3)}$ is not additive. For $|\psi\rangle^{AB}|\psi\rangle^{CD}$, $C_{(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\frac{1}{4}(h_{A}+h_{B}+h_{C}+h_{D})=\frac{1}{2}(h_{A}+h_{C})$ may be not larger than $C_{(2)}(AB|C|D)=2h_{C}/3$.
@@
$C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{h_{F}}/2<C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{DEF})=\sqrt{2h_{F}/3}$
@@
$C_{G,q(2)}(|\psi_{2}\rangle)=\sqrt{2h_{A}}/2<C_{G,q(2)}(|\psi\rangle^{AB}|\psi\rangle^{D})=\sqrt{2h_{A}/3},$
i.e., it is not coarsening monotonic. If $h_{C}>3h_{A}>0$, $C_{G,q(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\sqrt{(h_{A}+h_{C})/2}<C_{G,q(2)}(AB|C|D)=\sqrt{2h_{C}/3}$. So it is not tightly coarsening monotonic either. $C_{G,\alpha(k)}$ has the same property as that of $C_{G,q(k)}$.
@@ 5 Examples
We illustrate $E_{(k)}$, $\mathcal{E}_{(k)}$, and $E^{\prime}_{(k)}$ with the reduced functions $h_{C}(\rho)=\sqrt{2(1-{\rm Tr}\rho^{2})}$ and $h(\rho)=S(\rho)$, respectively. We denote them by $\acute{C}_{(k)}$ (in order to distinguish it from $C_{(k)}$ in Eq. (3)), $\mathcal{C}_{(k)}$, and $C^{\prime}_{(k)}$ if the reduced function is $h_{C}$, and by $E_{(k)}$, $\mathcal{E}_{(k)}$, and $E^{\prime}_{(k)}$ whenever the reduced function is $S$. Since $S$ and $h_{C}$ are subadditive Guo2024rip ; Wehrl1978 , so $E_{(k)}$, $\mathcal{E}_{(k)}$, $\acute{C}_{(k)}$, and $\mathcal{C}_{(k)}$ are complete. $E^{\prime}_{(k)}$ and $C^{\prime}_{(k)}$ are unified while $E^{\prime}_{(2)}$ and $C^{\prime}_{(2)}$ are complete.
@@ 6 Conclusion
We have established the axiomatic postulates for complete measure of the $k$-partite entanglement and presented two classes of $k$-partite entanglement measures. Comparing with the axiomatic postulates of complete $k$-entanglement measure in Ref. Guo2024pra , both the unification condition and the hierarchy condition were modified accordingly in order to make them in consistence with the type of entanglement considered. Together with the complete $k$-entanglement measure, we get a further progress in characterizing of multipartite entanglement. In comparison, although the $k$-PEM is far different from the $k$-entanglement measure, it has some similarities to the $k$-entanglement measure: all of them can be defined by the reduced function and in such a sense the completeness is always related to the subadditivity of the reduced function. In addition, we can discuss the monogamy and the complete monogamy relations of the $k$-entanglement measure, but it seems not compatible for $k$-PEM. Going further, our result is applicable for other $k$-partite measure of quantum correlations since it is based on the coarsening relation of partitions.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China under Grant Nos. 12471434 and 11971277, the Program for Young Talents of Science and Technology in Universities of Inner Mongolia Autonomous Region under Grant No. NJYT25010, and the High-Level Talent Research Start-up Fund of Inner Mongolia University under Grant No. 10000-A23207007.